OpenStudy (anonymous):

change to interval form and graph pleaseee! -21≤5x-9≤1

OpenStudy (precal):

add 9 to all 3 sides then divide all by 5

OpenStudy (anonymous):

\[-21\le5x-9\le1\]\[(+9)-21\le5x-9+9\le(+9)+1\]\[-12\le5x \le10\]\[\frac{ -12 }{ 5 }\le \frac{ 5x }{ 5 }\le \frac{ 10 }{ 5 }\]\[\frac{ -12 }{ 5 }\le x \le2\]